OpenStudy (anonymous):

Solve. 8 = 2x + 4 I know it's a positive number. My answer is 0

OpenStudy (anonymous):

8-4=4

OpenStudy (anonymous):

2(x+4)=8 x+4 = 8/2 = 4 x=4-4 = 0 yuuup ur right!!:)

OpenStudy (anonymous):

I got four at first, but then worked further along!!!!

OpenStudy (anonymous):

then divide by 2

OpenStudy (anonymous):

8=2x+4 4=2x x=2

OpenStudy (anonymous):

no x=2

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

x = -4 x = -1 x = 0 x = 7 ^Options

OpenStudy (anonymous):

its 0.. u were right:)

OpenStudy (anonymous):

thanks!!!! <333

OpenStudy (anonymous):

x is equal to 2...

OpenStudy (anonymous):

how r u guys getting 2?

OpenStudy (anonymous):

2 isnt an option....................

OpenStudy (anonymous):

Then are you sure the question isn't 8 = 2(x+4)?

OpenStudy (anonymous):

the answer is 0...